# Generalizable 3D Scene Reconstruction via Divide and Conquer from a Single View

### v1

@@ 3.3 Background modeling
Compared to the objects contained in a scene, the entities categorised as stuff, e.g. walls or ceiling, usually have a simpler geometry, which can be partially approximated by the corresponding regions in the depth map. Nonetheless, large areas are occluded by the foreground objects. Though we could follow a similar amodal completion strategy as the one described in 3.2, we opt for a more straightforward approach that interpolates the known data points. We consider all the background instances as one described by their mask union $M_{bg}$ and the corresponding image $I_{bg}$ and depth data $D_{bg}$. Then, we fit one small Multi-layer Perceptron $f:\mathbb{R}^{3}\to\mathbb{R}$ to represent the signed distance function (SDF) to the unprojected background points $P_{bg}^{view}$, and another one $c:\mathbb{R}^{3}\to\mathbb{R}^{3}$ to model the color associated to them. Lastly, we sample a dense grid of points in the camera frustum and evaluate their SDF values. The background surface is defined as $\mathcal{S}=\{x\in\mathbb{R}^{3}|f(x)=0\}$ and can be extracted from the grid using marching cubes [41].
@@ 4.1 Implementation details
As the proposed framework is not constrained to specific modules, we leverage the significant progress made by the computer vision community in the recent years towards solving the different sub-tasks which we describe in the method section.
In the scene analysis stage, we rely on CropFormer [53] for entity segmentation, OneFormer [24] for separating them in foreground instances and background entities, and Perspective Fields [26] for estimating the camera calibration. We mainly use Marigold [27] for depth estimation. However, the model predicts affine-invariant depth which differs by an unknown image-level offset and scale from the absolute physical units. During evaluations, we estimate these factors based on the ground truth depth available in the datasets to ensure that the reconstructions align with the target. For in-the-wild predictions, we empirically found that estimating the two unknowns of Marigold output based on a metric depth estimation, in our case, DepthAnything [66], achieves better results than using the latter by itself.
In the instance processing stage, we perform amodal completion on the reprojected crops using our model obtained by fine-tuning Stable Diffusion v1.5 [53]. We also sample several points in the instance’s mask and feed them to OVSAM[68] together with the input image to obtain the text prompt for guiding the diffusion. The completed object is then reconstructed using DreamGaussian[58]. As the model performs best with known camera elevation of the input view, we estimate it following the method introduced in [39]. Then, we find the 3D points of the reconstruction that correspond to the unprojected instance points, which serve as our layout guide, and compute the scale which aligns them.
@@ 4.2 Results
Table 1: As opposed to existing approaches, our method is at the same time compositional, able to generalize in a zero-shot manner, and does not require 3D-supervision for training.
Input image BUOL [9] Uni-3D [71] InstPIFu [37] Ours
Figure 4: Qualitative results on the 3D-FRONT dataset [15]. The methods consider reconstruct full scenes including background regions and foreground instances.
@@
Input image InstPIFu [37] DreamG[58] Ours GT geometry
Figure 5: Qualitative results on the 3D-FRONT [15] dataset illustrated under the input view and a second one chosen to highlight the reconstruction performance. For completeness, we also include the background in the original view, where it is available.
@@
DreamGaussian [58] 0.207 41.09 3.038 29.57
DreamGaussian
w/ our reprojection
@@
Table 2: Quantitative results on the reconstruction of foreground instances. On 3D-FRONT, we obtain comparable scores to InstPIFu which is trained supervised with 3D data on this dataset. No scores are reported for InstPIFu on the HOPE-Image dataset, as the method cannot recognize and reconstruct any of the objects included in the evaluation. On both datasets, we outperform the non-compositional approach, DreamGaussian.
We showcase the performance of our method using several datasets across diverse scenarios. For numerical evaluation we first consider 3D-FRONT [15, 16], a synthetic dataset of indoor rooms with available ground truth geometry. Due to the large scale of dataset, we manually sample 100 images from the test split of [37], avoiding the images with heavy scene occlusions (e.g., camera positioned behind a plant), unrealistic object intersections and scenes with very few objects. In addition, we use the 10 validation images of HOPE-Image [61] dataset containing household objects captured under two scenarios. Though the dataset is intended for object pose evaluation, we find the ground-truth object alignment to match the input images well-enough for our purpose.
We report the quantitative results using the widely-employed metrics for 3D reconstruction: Chamfer Distance [3] and F-Score [32]. Both are computed between densely sampled sets of points from the reconstructed meshes and the ground truth respectively. As we focus the evaluation on whole scenes, the points are uniformly sampled from the entire geometry of a scene. Therefore, larger objects impact the overall score more.
We compare our method against several solutions covering multiple approaches for 3D scene reconstruction. An overview of their capabilities is presented in table 1. BUOL[9] and Uni-3D [71] are both feed-forward scene reconstruction methods that have been trained with 3D supervision on the 3D-FRONT dataset. However, their training views are rendered in a simpler way compared to the renderings we use in our evaluation. Therefore, the poor performance of these methods evaluated numerically in table 3 and qualitatively in figure 4 shows their lack of generalization.
@@
InstPIFu [37] and USL [17] are both compositional approaches that rely on 2D and 3D object detectors for identifying the objects to be reconstructed and aligning them in the scene. This limits their reconstruction capabilities to a small number of classes. Being trained on similar scenes, InstPIFu has a significant advantage in our evaluation on the 3D-FRONT dataset. Nevertheless, our method surpasses its quantitative performance on full scene reconstruction (table 3) and closely matches the one on foreground reconstruction (table 2). The qualitative results in figures 4 and 5 show that the method successfully reconstructs large furniture pieces and arranges them in a good layout. However, several objects such as plants and chandeliers are disregarded. As the implementation of USL [17] is not publicly available, we only compare our visual results on the Hypersim dataset [52] in figure 7. Similar to InstPIFu, USL misses many objects, and the reconstructed components have a simplified geometry with no texture. In contrast, our results are much more realistic and have higher visual quality.
Additionally, while we use DreamGaussian [58] as part of our framework, we also compare with using DreamGaussian by itself in a non-compositional approach. To apply the model, we treat all the instances in a scene as a single object and reconstruct them together. As the model has seen several scenes composed of more than one object during its training [11], it performs reasonably under this setting. Given the different camera intrinsics in the evaluation, we also compare the results of applying the model on the images after using a reprojection similar to the one used in our pipeline. This further boosts its performance as measured in table 2. Still, its results are of significantly lower quality compared to our compositional approach, which can be analyzed in figure 6.
Lastly, we qualitatively evaluate in figure 8 the performance of InstPIFu [37], DreamGaussian [58] and our method on real-world data with diverse scenarios. The results show that the proposed solution reconstructs complex scenes well, while overcoming many limitations of prior works.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Input image DreamGaussian[58] Ours Ground Truth
Figure 6: Qualitative results on HOPE-Image [61].
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Figure 8: Qualitative results on real-world data.
@@ 4.3 Ablations
The two most important interfaces in combining the depth estimation and the single-view object reconstruction components are our reprojection and amodal completion. We present an ablation of these components in table 4. The object reconstruction module $\mathcal{R}$ expects an image of an object centered in the camera frustum. Therefore, simply feeding an image crop from different parts of the scene image will lead to unrealistic views and, consequently, deformed output objects. Additionally, we use amodal completion to contend with the various occlusions that appear in the input view. This step significantly boosts the performance on the 3D-FRONT dataset, but does not improve the numerical evaluation on HOPE-Image dataset, where most of the objects in the scenes are not occluded.
@@
Table 4: Ablation results for our reprojection and amodal completion procedures. Both components contribute to the performance of our method.
@@ 4.4 Limitations
The proposed method has certain shortcomings and there is significant room for improvement, especially for in-the-wild predictions. As we rely on the unprojected depth for the layout of the scene, when the camera calibration or the scale and offset estimation fail to match the correct units, the reconstructed scene will have an erroneous structure, with objects colliding with one another.
We also inherit the limitations of the single-view object reconstruction method. In our experiments, we notice that DreamGaussian sometimes overestimates the size of the object along the $z$-dimension, when it cannot be observed in the input view. Such cases are prominent in figure 6, e.g., orange juice box in the last example. Additionally, DreamGaussian requires the estimation of the perceived camera elevation of the input crop with respect to the normalized object pose. As this cannot be robustly inferred, it significantly affects the reconstruction performance of the model. We believe that most of the current limitations can be overcome by improving the implementation of some of the particular modules in our framework and by enhancing their interoperability. For example, conditioning the depth estimation on the inferred camera calibration can help disambiguate depth scale [55]. Also, the 3D object reconstruction module can benefit from having the estimated instance depth as an additional input [22, 62]. Lastly, relying more on the overall scene context during the amodal completion stage can improve the recovery of missing object parts due to occlusion, as it has been recently shown in concurrent works [46, 64].
@@ 5 Conclusion
In this work, we introduce a modular framework for reconstructing complex 3D scenes from an image. We prioritize generalization by taking a divide-and-conquer approach as opposed to an end-to-end model. Our decomposing into multiple entities benefits from existing components, which effectively solve the established subtasks. We develop the necessary interfaces that enable the sub-components to function properly and finally yield a full 3D reconstruction. Our experiments decidedly show the advantage of the proposed method on various types of scenes. Considering the illustrated performance on the diverse scenarios, we believe that the proposed method is a strong baseline and a stepping stone towards achieving generalizable full 3D scene reconstruction from a single image.
Acknowledgement
@@
This work was funded by the German Federal Ministry of Education and Research (BMBF), FKZ: 01IS22082 (IRRW). The authors are responsible for the content of this publication. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b112dc IRRW. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683.

### v2

@@ 3.3 Background modeling
Compared to the objects contained in a scene, the entities categorised as stuff, e.g. walls or ceiling, usually have a simpler geometry, which can be partially approximated by the corresponding regions in the depth map. Nonetheless, large areas are occluded by the foreground objects. We consider all the background instances as one described by their mask union $M_{bg}$ and the corresponding image $I_{bg}$ and depth data $D_{bg}$. Then, we fit one small Multi-layer Perceptron $f:\mathbb{R}^{3}\to\mathbb{R}$ to represent the signed distance function (SDF) to the unprojected background points $P_{bg}^{view}$, and another one $c:\mathbb{R}^{3}\to\mathbb{R}^{3}$ to model the color associated to them. Lastly, we sample a dense grid of points in the camera frustum and evaluate their SDF values. The background surface is defined as $\mathcal{S}=\{x\in\mathbb{R}^{3}|f(x)=0\}$ and can be extracted from the grid using marching cubes [49, 70].
@@ 4 Experiments
Input image BUOL [11] Uni-3D [85] InstPIFu [45] Ours
Figure 4: Qualitative results on the 3D-FRONT dataset [19]. The methods considered reconstruct full scenes including background regions and foreground instances.
@@ 4.1 Implementation details
As the proposed framework is not constrained to specific modules, we leverage the significant progress made by the computer vision community in the recent years towards solving the different sub-tasks described above.
In the scene analysis stage, we rely on CropFormer [59] for entity segmentation, OneFormer [31] to separate them in foreground instances and background entities, and Perspective Fields [33] to estimate the camera calibration. We mainly use Marigold [34] for depth estimation. However, the model predicts affine-invariant depth which differs by an unknown image-level offset and scale from the absolute physical units. During evaluations, we estimate these factors based on the ground truth depth available in the datasets to ensure that the reconstructions align with the target. For in-the-wild predictions, we empirically found that estimating the two unknowns of Marigold output based on a metric depth estimation, in our case, DepthAnything [80], achieves better results than using the latter by itself.
In the instance processing stage, we perform amodal completion on the reprojected crops using our model obtained by fine-tuning Stable Diffusion v1.5 [64]. We also sample several points in the instance’s mask and feed them to OVSAM [82] together with the input image to obtain the text prompt for guiding the diffusion. The completed object is then reconstructed using DreamGaussian [69]. We estimate the camera elevation required by DreamGaussian as in [46]. Then, we find the 3D points of the reconstruction that correspond to the unprojected instance points, which serve as our layout guide, and compute the scale which aligns them. Further specifications regarding the integrated models, possible alternatives for some of the processing stages and an analysis of the inference time of our method are provided in Section 6 of the supplementary material.
@@ 4.2 Results
Table 1: As opposed to existing approaches, our method is at the same time compositional, able to generalize in a zero-shot manner, and does not require training with 3D supervision.
We showcase the performance of our method using several datasets across diverse scenarios. For numerical evaluation we first consider 3D-FRONT [19, 20], a synthetic dataset of indoor rooms with available ground truth geometry. Due to the large scale of the dataset, we manually sample 100 images from the test split of [45], avoiding the images with heavy scene occlusions (e.g., camera positioned behind a plant), intersecting objects, and scenes with very few objects. In addition, we use the 10 validation images of HOPE-Image [73] dataset containing household objects captured under two scenarios. Though the dataset is intended for object pose evaluation, we find the ground-truth object alignment to match the input images well-enough for our purpose.
We report the quantitative results using the widely-employed metrics for 3D reconstruction: Chamfer Distance [4] and F-Score [39]. Both are computed between densely sampled sets of points from the reconstructed meshes and the ground truth respectively. As we focus the evaluation on whole scenes, the points are uniformly sampled from the entire geometry of a scene.
We compare our method against several solutions covering multiple approaches for 3D scene reconstruction. An overview of their capabilities is presented in Table 1. BUOL [11] and Uni-3D [85] are both feed-forward scene reconstruction methods that have been trained with 3D supervision on the 3D-FRONT dataset. While we evaluate the methods on the same dataset they were trained on, we use a more realistic rendering, following [45]. Even under this minor change, the methods’ performance degrades significantly, as can be seen in Table 3 and Figure 4, showing their lack of generalization.
@@
Input image InstPIFu [45] DreamGaussian [69] Ours GT geometry
Figure 5: Qualitative results on 3D-FRONT [19] illustrated under the input view and a second one chosen to highlight the reconstruction performance. In contrast to Ours and InstPIFu, DreamGaussian reconstructs all the objects at once and does not model the background.
@@
DreamG [69] 0.207 41.09 3.038 29.57
DreamG $+$
reprojection
@@
Table 2: Quantitative results on foreground instances reconstruction. On 3D-FRONT, we are on par with InstPIFu, despite it being trained with 3D data on this dataset. On HOPE-Image, InstPIFu fails to recognize any of the evaluated objects. On both datasets, we outperform the non-compositional approach, DreamGaussian.
@@
InstPIFu [45] and USL [22] are both compositional approaches that rely on 2D and 3D object detectors for identifying the objects to be reconstructed and aligning them in the scene. However, InstPIFu requires direct 3D supervision and USL is trained end-to-end. This limits their application domain and generalizability. Furthermore, as seen in Tables 3 and 2, our method is able to quantitatively match the performance of InstPIFu even when evaluated on the 3D-FRONT dataset (which InstPIFu used for training). The qualitative results in Figures 4 and 5 show that the method successfully reconstructs large furniture pieces and arranges them in a good layout; however, several objects such as plants and chandeliers are missing. The results of InstPIFu further degrade when evaluated out-of-distribution, as can be seen on real-world images in Figure 7. Since the implementation of USL [22] is not publicly available, we only compare our visual results on the Hypersim dataset [63] in Figure 6. USL also misses many objects, and the reconstructed components have a simplified geometry with no texture. In contrast, our results are more realistic and have higher visual quality.
Additionally, we compare our method with DreamGaussian [69] by itself in a non-compositional approach. To apply the model, we treat all instances in a scene as a single object and reconstruct them together. As the model has seen several scenes composed of more than one object during its training [14], it performs reasonably under this setting. Given the different camera intrinsics in the evaluation, we also compare the results of applying the model on the images after using a reprojection similar to the one used in our pipeline. This further boosts its performance as measured in Table 2. Still, its results are worse compared to our compositional approach, which can be analyzed in the qualitative results in Figure 5 and in the supplementary material.
Lastly, we evaluate in Figure 7 the methods’ performance on real-world data with diverse scenarios. The results show that the proposed solution reconstructs complex scenes well, while overcoming many limitations of prior works. We briefly address the reconstruction of outdoor scenes in the Section 7.4 of the supplementary material.
@@
Figure 7: Qualitative results on real-world data. Interactive visualizations are included on our project page.
@@ 4.3 Ablations
The two most important interfaces in combining the depth estimation and the single-view object reconstruction components are our reprojection and amodal completion. We present an ablation of these components in Table 4. Ablating the reprojection amounts to simply using image crops. This ignores projective geometry properties and leads to deformed reconstructions. Amodal completion is necessary to contend with the various occlusions that appear in the input view. This step boosts the performance on the 3D-FRONT dataset, but does not improve the numerical evaluation on HOPE-Image dataset, since most of the objects in the scenes are not occluded. More ablation results are included in the Section 7.3 of the supplementary material.
@@
w/o reprojection 0.155 58.56 1.505 51.89
w/o amodal comp. 0.122 66.81 1.450 54.99
Table 4: Ablation results for our reprojection and amodal completion modules; both contribute to the performance of our method.
@@ 4.4 Limitations
The proposed method has certain shortcomings and there is significant room for improvement, especially for in-the-wild predictions. By design, the failure cases of the individual modules (depth estimation, camera calibration, elevation estimation, etc.) become limitations of our framework. Since we do not train an end-to-end system, errors can propagate from one stage to the next. Therefore, the performance of the overall pipeline is limited by its weakest link.
We believe that most of the current limitations can be overcome by improving the implementation of some of the particular modules in our framework and by enhancing their interoperability: using estimated depth in 3D object reconstruction or global image context for amodal completion. We further discuss the method’s limitations and provide concrete examples in the Section 8 of the supplementary material.
@@ 5 Conclusion
In this work, we introduce a modular framework for reconstructing complex 3D scenes from an image. We prioritize generalization by taking a divide-and-conquer approach rather than end-to-end. Our decomposing into multiple entities benefits from existing components, which effectively solve the established subtasks. We develop the necessary interfaces that enable the modules to function properly and finally yield a full 3D reconstruction. Our experiments decidedly show the advantage of the proposed method on various types of scenes. Considering the illustrated performance for diverse scenarios, we believe that our approach is a strong baseline and a stepping stone towards generalizable full 3D scene reconstruction from a single image.
@@
Acknowledgement. This work was funded by the German Federal Ministry of Education and Research (BMBF), FKZ: 01IS22082 (IRRW). The authors are responsible for the content of this publication. The authors appreciate the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b112dc IRRW. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683.
Supplementary Material
Overview
This supplementary material includes additional details and results which provide insights into our novel approach for generalizable 3D scene reconstruction. In section 6, we present supporting information for the reproducibility of our method and the evaluation setup. Section 7 extends the ablation study in the main paper and shows the effect of replacing different modules in the pipeline (7.2), presents a qualitative comparison for the ablations (7.3), and discusses the application on outdoor scenes (7.4). Then, in section 8, we elaborate and illustrate the limitations of our method. Additional results and our code base can be found on the project page  [^1] https://andreeadogaru.github.io/Gen3DSR.
6 Implementation details
6.1 Amodal completion
Amodal completion differs from inpainting by ensuring that the model does not hallucinate new objects when recovering the missing parts, as shown in Figure 9.
The amodal completion step in our framework is fulfilled by an image-to-image Stable Diffusion [64] model fine-tuned on a custom dataset. We create this dataset starting from OmniObject3D [77] which contains $6000$ high-quality real-scanned 3D objects. Though the model only needs 2D images for training, we found it beneficial to render 3D objects instead of using real images. This allows us to obtain well-segmented singular objects. The Blender-rendered images of the objects are used as the target images. To obtain the conditioning view, we mask-out parts of the object by randomly overlaying the silhouette of an arbitrary object sampled from the Objaverse dataset [13]. All the channels of the background and the occluded pixels are set to the same value of $127.5$ (equivalent to zero after image normalization). The category labels provided in OmniObject3D are used as prompts to guide the diffusion process. Training samples from the dataset are provided in Figure 8. We train the model at a resolution of $512\times 512$ for $25000$ iterations with a batch size of $16$.
As our amodal completion model does not explicitly infer the amodal mask of the completed object, we obtain it by using a foreground segmentation model [61].
Figure 8: Training samples from the amodal completion synthetic dataset with objects from OmniObject3D [77].
Input crop SD XL inpainting Our amodal completion
Figure 9: In comparison to our amodal completion, inpainting models tend to hallucinate new objects when filling in the holes. Both models use desk as prompt.
6.2 Reprojection and view-space alignment
Utilizing the estimated depth map $D$ alongside the camera calibration $K_{img}$, we perform pixel unprojection from the input image, resulting a point cloud $P^{view}$ within the view space. We refer to $P^{view}$ as our layout guide; this serves as a pivotal reference, ensuring the alignment of all individually reconstructed components within the view space, thus forming the complete scene. While the background modeling step directly fits a SDF to the corresponding background points $P_{bg}^{view}$, the instance processing step uses $P_{i}^{view}$ for two purposes: reprojection and alignment.
The straightforward approach of simply cropping the instances out of the original image and feeding them to the single-shot object reconstruction method $\mathcal{R}$ leads to deformed reconstructions. This happens because the view-conditioned diffusion model $\mathcal{Z}$, which is used for generating novel views of the object, assumes images are captured under a predefined setup [48]. Specifically, the object should be in the center of the image, captured by a camera with field of view of $49.1^{\circ}$, positioned at a distance between $[1.0,1.7]$ from the object which is normalized to fit into the unit cube. As these conditions are not satisfied by arbitrary crops within the image, we project $P_{i}^{view}$ into a crop $C_{i}$ with similar properties. The virtual camera employed for projection uses the $\mathcal{Z}$-compatible camera setup; i.e., a camera positioned at a distance of $1.5$ from the normalized $P_{i}^{view}$ and oriented towards its center, with intrinsics $K_{crop}$. The reconstructed object might have a different scale compared to $P_{i}^{view}$. Therefore we estimate the scale factor $s_{i}$ that aligns the reconstructed object with $P_{i}^{view}$ using a RANSAC-based approach [18], accounting for the possible mismatches.
6.3 Background modeling
The MLP used to reconstruct the background has 4 hidden layers with 128 neurons each and Softplus activations. The network is trained from scratch for each scene using points sampled along background camera rays. The SDF of the points is computed as the distance to the unprojected estimated depth. As we do not sample points in the regions originally occluded, the network implicitly interpolates the missing areas based on the visible surrounding regions.
6.4 Integrated models
Our method ensembles several models, each tackling a different sub-tasks as follows:
  • Camera calibration (estimation of field of view and principal point): Perspective Fields [33] trained on 360cities [1] and EDINA [15] datasets.
  • Entity segmentation: CropFormer [60] with Hornet-L backbone trained on EntitySeg dataset [60].
  • stuff-thing segmentation: OneFormer [31] with DiNAT-L backbone trained on ADE20K [86]. The model predicts 150 classes which are grouped in stuff and thing. We make some modifications to the original grouping; thing $\to$ stuff: window, door, curtain, mirror, fence, rail, column, stairs, screen door, balustrade, step; stuff $\to$ thing: plant, tent, crt screen, cradle, blanket.
  • Monocular depth estimation: Depth Anything [80] fine-tuned on NYUv2 [54] (metric depth) and Marigold [34] with Stable Diffusion v2 backbone trained on two synthetic datasets (affine-invariant depth).
  • Object recognition: OVSAM [82] which combines SAM [38] with CLIP[32] and is trained on COCO [44] and LVIS [23] datasets. For each object we sample 5 points inside the eroded instance mask to prompt the model.
  • Amodal completion: our model based on Stable Diffusion v1.5 and trained on the synthethic dataset described in section 6.1 of this supplementary material.
  • Singe-image object reconstruction: DreamGaussian [69] which reconstructs the object using 3D Gaussians [35] and employs Zero-1-to-3 XL [48] trained on ObjaverseXL [14] as the 2D diffusion prior.
We show variants of our pipeline with different modules for some of the processing steps (depth estimation, stuff-thing segmentation, and object reconstruction) in section 7.2.
6.5 Inference time
Using the default pipeline configuration on an NVIDIA RTX A5000, for an image with resolution $1500\times 1500$, the scene analysis stage takes $\approx 80$ sec and for each instance specific processing additional $\approx 90$ sec are required. Similar to other compositional methods (e.g., InstPIFu [45]), the inference time increases linearly with the complexity of the scene (number of objects). This limitation can be mitigated by paralellizing the reconstruction of the objects, provided the available hardware resources allow it. Still, the time per instance is quite high, mostly due to the inference time of DreamGaussian [69]. This particular method optimizes 3D Gaussians to views of the object and has an additional step for texture refinement, adding to the overall runtime. However, as we designed a modular pipeline, we can replace DreamGaussian [69] with new, improved methods for single-view 3D object reconstruction, such as LaRa [8]. This very recent method enables us to reduce the processing time per instance to $\approx 50$ sec. Finally, the processing time could be further improved by streamlining the pipeline for a specific configuration of modules.
6.6 Evaluation details
For 3D-FRONT [19, 20] we report the metrics at the original scale of the geometry in the dataset using 1000000 points sampled from both the reconstructed meshes and the ground truth geometry. The F-Score is computed at a threshold of $0.1$. As most of the modules in our pipeline operate at a high-resolution, the original resolution of the images ($648\times 484$) is insufficient. Therefore, we first increase their size to $1296\times 968$ using a super-resolution method [76].
For HOPE-Image [73], we compose the ground truth geometry using the provided pose annotations. Additionally, we found that a scale of $0.1$ must first be applied to the original object meshes to match with the images. The metrics are reported at this scale based on 500000 sampled points and the threshold used for F-Score is $1.0$. On this dataset, we only evaluate the reconstruction of the foreground instances, as there is no ground truth background geometry.
7 Additional experiments
7.1 HOPE-Image
Input image DreamGaussian [69] Ours Ground Truth
Figure 10: Qualitative results on HOPE-Image [73]. We show each reconstruction from two camera views, including the input one.
We include in Figure 10 the qualitative results of the experiments on HOPE-Image dataset [73]. We compare our compositional approach with DreamGaussian [69], which reconstructs all the objects in the image at once. Our method can better handle complex scenes with many objects, as each instance is individually reconstructed. In contrast, both the appearance and the geometry of DreamGaussian’s reconstructions degrades when applied on scenes with multiple objects. This is mainly due to limitations of the Zero-1-to-3 XL method, which fails to generate realistic, multi-view-consistent views. The pitfall is expected, as the domain required to be modeled by the prior increases exponentially with the number of objects. We do not include InstPIFu [45] in the comparison on HOPE-Image dataset because the method fails to detect any of the considered objects in the scenes.
7.2 Alternative models
Depth estimation stuff vs thing Object reconstruction Chamfer $\downarrow$ F-Score $\uparrow$
Marigold OneFormer DreamGaussian 0.099 75.33
[…]
Marigold CLIPSeg DreamGaussian 0.166 65.56
Marigold OneFormer One-2-3-45 0.110 69.40
Table 5: Quantitative evaluation of our method using alternative models on 3D-FRONT [19] dataset.
The proposed method is designed as a modular framework, allowing the straightforward replacement of the integrated models summarized in 6.4. We showcase this property of our method by exchanging the Marigold [34] depth estimation with Depth Anything [80], the stuff-thing detection based on OneFormer [31] with one based on CLIPSeg [50], and the single-view object reconstruction model, DreamGaussian [69], with One-2-3-45 [46]. We evaluate in Table 5 the performance of these modifications on full scene reconstruction (background and foreground objects) using the 3D-FRONT [19] dataset. For CLIPSeg, we define empirically a set of prompts for foreground (object, furniture) and background (background, floor, wall, curtain, window, ceiling), and consider foreground pixels the ones that have a lower than $0.5$ score for all the background prompts or higher than $0.1$ for any of the foreground ones. The performance of our method decreases in this case because DreamGaussian poorly reconstructs the background regions that are misclassified as things. We observe this happens more frequently on indoor rooms when using CLIPSeg, and on tabletop scenes, such as the one in our method figure from the main paper, when using OneFormer.
The results show that our default configuration (as reported in the main paper) yields the best metrics on this dataset. Still, the alternative configurations are competitive, which proves that our pipeline is flexible and robust to different implementations of the modules.
7.3 Ablations
We provide in Figure 11 qualitative results corresponding to the two ablations considered in the main paper (see section 4.3). When ablating the amodal completion step, the reconstructed objects have holes or incorrect textures, corresponding to the regions that were occluded by other objects in the scenes. When skipping the reprojection step of the pipeline and directly using crops from the input image, the objects are reconstructed with deformed shapes to match the input view. By including both components, our full method reconstructs complete objects with a physically-correct geometry. We also consider these ablations together with LaRa [8], a very recent method for single view object reconstruction that operates in a feed-forward fashion. Compared to DreamGaussian which relies on Zero-1-to-3 XL to generate novel views conditioned on the input crop, LaRa uses a newer model, Zero123++, that can generate six fixed views that are more 3D consistent. As can be seen in Figure 12, the proposed amodal completion and reprojection are also beneficial for LaRa, boosting the reconstruction performance.
OursDG
w/o reprojection
Refer to caption Refer to caption Refer to caption
OursDG
w/o amodal
completion
Refer to caption Refer to caption Refer to caption
OursDG
Refer to caption Refer to caption Refer to caption
OursDG
w/o reprojection
Refer to caption Refer to caption Refer to caption
OursDG
w/o amodal
completion
Refer to caption Refer to caption Refer to caption
OursDG
Refer to caption Refer to caption Refer to caption
Method Input crop Reconstructed object
Figure 11: Ablations visual comparison. In OursDG we use DreamGaussian [69] for reconstructing individual objects with the proposed reprojection of the input pixels and amodal completion of the missing parts.
OursLaRa
w/o reprojection
Refer to caption Refer to caption Refer to caption
OursLaRa
w/o amodal
completion
Refer to caption Refer to caption Refer to caption
OursLaRa
Refer to caption Refer to caption Refer to caption
OursLaRa
w/o reprojection
Refer to caption Refer to caption Refer to caption
OursLaRa
w/o amodal
completion
Refer to caption Refer to caption Refer to caption
OursLaRa
Refer to caption Refer to caption Refer to caption
Method Input crop Reconstructed object
Figure 12: Ablations visual comparison. In OursLaRa we use LaRa [8] for reconstructing individual objects with the proposed reprojection of the input pixels and amodal completion of the missing parts.
7.4 Outdoor scenes
Our method can reconstruct well a wide range of scenes, from tabletop setups to large rooms with many objects of varying sizes. Though we exemplify the performance of the proposed pipeline on indoor environments, by design the method is also applicable to outdoor scene reconstruction. In Figure 13, we provide qualitative results for such scenes. In these experiments, we use Marigold [34] as a depth estimator and compute the scale and shift based on Depth Anything [80] prediction of the model fine-tuned on KITTI [21] (suited for outdoor scenes).
We choose to focus more on indoor scene reconstruction because most outdoor scenes are predominantly composed of entities categorized as stuff, which we represent simply with a mesh approximating the estimated depth map. We believe a dedicated approach for outdoor background reconstruction would be better suited for this type of scenes. Nonetheless, outdoor scenes with several objects are reconstructed reasonably well by our method, as can be observed in the qualitative results.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Input image Ours
Figure 13: Qualitative results of our method on real-world outdoor scene reconstruction.
8 Limitations
We describe in this section some of the limitations our method inherits from the integrated modules.
As we rely on the unprojected depth for the layout of the scene, when the camera calibration or the scale and offset estimation fail to match the correct units, the reconstructed scene will have an erroneous structure. We provide an example of incorrect camera calibration in Figure 15. Furthermore, since entities are reconstructed independently, without relative constraints between the objects, the method does not address potential physical impossibilities such as intersecting objects.
In our experiments, we notice that DreamGaussian sometimes overestimates the size of the object along the $z$-dimension, when it cannot be observed in the input view. Such cases are prominent in Figure 10, e.g., orange juice box in the last example. Additionally, DreamGaussian requires the estimation of the perceived camera elevation of the input crop with respect to the normalized object pose. Since this cannot be robustly inferred, it significantly affects the reconstruction performance of the model, as can be analysed in Figure 14.
Though our approach for modeling the background is superior to previous works, using a small MLP to model the ‘stuff’ of the scene (entities that are not objects) may sacrifice details in the texture. Also, the implicit interpolation for the occluded areas is sometimes insufficient to cover large missing regions, resulting in holes in the background surface.
As the proposed method follows a modular approach, we can improve on this limitations by upgrading the particular modules. For example, conditioning the depth estimation on the inferred camera calibration can help disambiguate depth scale [66, 26]. Also, the 3D object reconstruction module can benefit from having the estimated instance depth as an additional input [29, 74]. Lastly, relying more on the overall scene context during the amodal completion stage can improve the recovery of missing object parts due to occlusion, as it has been recently shown in [56, 78].
Input view $-35^{\circ}$ $-30^{\circ}$ $-25^{\circ}$ $-20^{\circ}$ $-15^{\circ}$
Figure 14: DreamGaussian [69] object reconstructions of the input view under different camera elevation angles. We illustrate the 3D reconstructions using a camera view from the side to highlight their differences. The elevation angle estimated by the method proposed in [46] is $-35^{\circ}$. However, the best reconstruction is achieved using an elevation angle of around $-25^{\circ}$.
Refer to caption Refer to caption Refer to caption
Input image Predicted FOV Adjusted FOV
Figure 15: Comparison of unprojected depth under two camera settings. The field of view prediction of PerspectiveFields [33] is incorrect for this image, resulting in a flattened scene. The unprojected depth using a camera with a manually adjusted the field of view has more realistic proportions.